OpenStudy (lilai3):

Please help me-Algebra 1! WILL REWARD MEDAL/FAN!!! Fiona and Greg play a number guessing game. Greg gives Fiona this hint about his two secret numbers, "The product of the two consecutive positive integers that I am thinking of is 11 more than their sum." What are Greg's numbers? Please help.

OpenStudy (dan815):

x*(x+1)=11+(x+1)+x

OpenStudy (shrutipande9):

let the 2 nos be x and x+1 x(x+1)=11+x+x+1

OpenStudy (dan815):

simplify and solve

OpenStudy (lilai3):

Yes, I simplified and got the answer: